shift summary day shift Dec 5,1999

Mike Arenton (arenton@uvahea.phys.virginia.edu)
Sun, 5 Dec 1999 16:30:42 -0500 (EST)

Shift Summary for day shift, 12/05/99

On Shift: Mike Arenton

Hours of beam delivered at all intensities:0
Hours of beam delivered at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...):

Accounting of lost time:
TeV magnet replacement

===============================================================

Runs ended this shift: none

===============================================================

Tapes brought to FCC this shift: none

Accesses during shift: open access

===============================================================

Description of Shift:

Waiting while magnets get replaced. J.Shields running
trigger 1 skim's on ktev drives.

Rick K. does tests of software for KAMI test. Note that
he unplugged a cable (having to due with calibration events
to monitor plane) and next shift should verify that he
plugs it back in before he leaves.

First snowfall of season outside.

There are 2 dead bulbs in holiday lights.

Notes for next shift:

SA3 HV is still bad, the power supply still needs to be
replaced (probably). The HV comes back up if you unplug the
HV output cable to the cow. However there isn't any single channel
that's pulling it down, so it's probably the supply.